# SelectTTS: Synthesizing Anyone’s Voice via Discrete Unit-Based Frame Selection

### v1

@@ IV-B Vocoder
TABLE I: Objective evaluation results (arrows indicate whether higher or lower metric values are better)
IV-C Baselines
We use three SOTA multi-speaker TTS baselines to compare the performance of our proposed framework: the official releases of YourTTS [^2] https://github.com/Edresson/YourTTS [7] and XTTS-v2333https://huggingface.co/coqui/XTTS-v2 [3] with default parameters. Since there is no official release of VALL-E [2], we use Amphion’s implementation [^4] https://github.com/open-mmlab/Amphion/tree/main/egs/tts/VALLE_V2.
IV-D Training and test data
We utilize LibriSpeech train-clean-100 dataset [35] for training both the text-to-semantic-unit tokenizer and HiFi-GAN vocoder. This dataset comprises of 100 hours of text-audio pairs, sampled at 16kHz, and features 251 speakers. We use the LibriSpeech dev-clean subset for validation and LibriTTS-R-test-clean [36] for testing. For the test set, we select speakers from the subset who have at least 20 sentences, each with a word count between 3 and 30. This criteria results in 31 eligible speakers and a total of 620 utterances. Additionally, for the reference speech, we use 5 minutes of audio across all variants of SelectTTS, YourTTS, and XTTS-v2. However, for VALL-E, due to constraints of the language modeling, the reference speech was limited to a duration of 10-15 seconds.
V Results and discussion
We include four variants of the proposed model in our evaluations. The first variant (inv-kmeans (rand)) uses inverse k-means sampling with a randomly selected frame from the predicted class. The second variant (inv-kmeans (avg)) also uses inverse k-means sampling but averages all frames within the predicted class for decoding. The third and fourth variants build on the first two by incorporating sub-sequence matching.
Method #Parameters Training data (hours) RTF $\downarrow$
VALL-E 594M 45k 1.325
XTTS-v2 466M 27k 0.186
YourTTS 86M 474 0.120
SelectTTS w/o sub-match 57M 100 0.290
SelectTTS w sub-match 57M 100 0.334
TABLE II: Analysis of model complexity
@@ V Results and discussion
Ground Truth 4.36 $\pm$ 0.09 -
VALL-E 4.22 $\pm$ 0.08 3.74 $\pm$ 0.11
XTTS-v2 4.14 $\pm$ 0.09 3.77 $\pm$ 0.10
YourTTS 3.15 $\pm$ 0.10 3.25 $\pm$ 0.11
SelectTTS w/o sub-match 3.90 $\pm$ 0.09 3.85 $\pm$ 0.09
SelectTTS w sub-match 4.04 $\pm$ 0.09 3.82 $\pm$ 0.08
TABLE IV: Subjective evaluations (95% confidence interval)
V-A Objective evaluation
Table I presents the objective evaluation results, where we assess the intelligibility of the synthesized speech using WER (word error rate) calculated with the Wav2Vec 2.0 Large ASR model [^5] https://huggingface.co/facebook/wav2vec2-large-960h-lv60-self. Additionally, we measure naturalness with UTMOS [37], and observe that sub-sequence matching significantly improves the performance. These results indicate that our proposed method beats YourTTS in intelligibility and naturalness and comes close to much larger baselines like VALL-E and XTTS-v2. We highlight that SelectTTS outperforms all baselines in SECS (Speaker Encoder Cosine Similarity) computed using ECAPA-TDNN speaker embeddings[38]. Moreover, model complexity analysis in Table II shows that SelectTTS, with 8x fewer parameters than XTTS-v2 and 10x fewer than VALL-E, while requiring 270x and 450x less training data respectively, is significantly lighter while maintaining a comparable Real-Time Factor (RTF). The ablation study in Table III, demonstrates how the reference speech duration impacts WER and SECS performance. Even when the reference speech is short, SelectTTS remains competitive with 30 seconds of reference speech, despite the higher WER and lower SECS due to limited frames in the predicted cluster.
@@ V-B Subjective evaluation
We conduct listening experiments with 15 subjects to assess the naturalness and speaker similarity of the synthesized speech, using a total of 120 utterances from 6 unseen speakers (3 male, 3 female) for the tests. In the Mean Opinion Score (MOS) evaluation [7], participants rate the naturalness of the speech using a 5-point scale for 12 generated speech samples per method. As indicated in Table IV, the proposed SelectTTS generates speech with a high level of naturalness comparable with strong baselines, achieving an MOS score close to 4. We can also see the positive effect of sub-sequence matching algorithm in naturalness. We also conduct a Speaker Mean Opinion Score (SMOS) [7] in which participants are presented with the ground truth alongside the synthesized speech from the proposed method and baselines. They are then asked to rate the speaker similarity on a 5-point scale for 12 speech samples per method. As the results in Table IV indicate, our method has very strong SMOS with high speaker resemblance to the target speaker, better or on par with larger baselines.
@@ VI Conclusion
In this paper, we propose SelectTTS, a multi-speaker TTS framework with lower model complexity that directly utilizes frames from the unseen target speaker to synthesize high-quality speech that closely resembles the target speaker’s voice. We demonstrate that our approach of combining frame selection based on semantic units with vocoding using a sequence of SSL features offers a much simpler yet effective method for modeling unseen speakers in the multi-speaker TTS task attaining SOTA target speaker similarity.

### v2

@@ IV-C Training and test data
TABLE II: Objective evaluation results (VCTK)
Method WER(%$\downarrow$ SECS $\uparrow$ UTMOS $\uparrow$
Ground Truth 4.39 69.77 4.04
VALL-E 8.53 46.70 4.07
XTTS-v2 4.89 54.12 4.00
YourTTS 9.90 36.83 3.53
SelectTTS (only inv k-means (avg)) 7.18 60.61 3.69
SelectTTS (inv k-means (avg) + sub-match) 7.01 54.13 3.92
@@
V Results and discussion
We evaluate four variants of the proposed SelectTTS framework. The first variant, inv-kmeans (rand), employs inverse $k$-means sampling by randomly selecting a frame from the predicted cluster. The second variant, inv-kmeans (avg), also uses inverse $k$-means sampling but decodes using the average of all frames within the predicted cluster. The third and fourth variants build on the first two by incorporating sub-sequence matching.
V-A Objective evaluation
Table I presents objective evaluation results on the LibriTTS-R test set, measuring intelligibility via WER using the Wav2Vec 2.0 Large ASR model [^5] https://huggingface.co/facebook/wav2vec2-large-960h-lv60-self, and naturalness via UTMOS [35]. We observe that sub-sequence matching improves both metrics. SelectTTS outperforms YourTTS in intelligibility and naturalness, and approaches the performance of larger models like VALL-E and XTTS-v2. SelectTTS also achieves the highest speaker similarity across all models, as measured by SECS using ECAPA-TDNN embeddings [36]. Results on the VCTK test set (Table II) show a similar trend: the proposed method attains the best speaker similarity with competitive WER and UTMOS scores, demonstrating robustness across datasets. Table III analyzes the impact of reference speech duration. Although performance slightly drops with shorter references, SelectTTS remains competitive even with only 30 seconds of reference audio. As shown in Table IV, SelectTTS requires 8× fewer parameters than XTTS-v2 and 10× fewer than VALL-E, with 270× and 450× less training data, respectively, while maintaining a comparable Real-Time Factor (RTF).
TABLE IV: Analysis of model complexity
Method #Params Train data (hours) RTF $\downarrow$
VALL-E 594M 45k 1.325
XTTS-v2 466M 27k 0.186
YourTTS 86M 474 0.120
SelectTTS w/o sub-match (inv k-means (avg)) 57M 100 0.290
SelectTTS w sub-match (inv k-means (avg)) 57M 100 0.334
TABLE V: Subjective evaluations (95% confidence interval)
@@ V-A Objective evaluation
Ground Truth 4.20 $\pm$ 0.09 -
VALL-E 4.11 $\pm$ 0.08 3.62 $\pm$ 0.09
XTTS-v2 4.10 $\pm$ 0.08 3.64 $\pm$ 0.10
YourTTS 3.15 $\pm$ 0.10 3.22 $\pm$ 0.09
SelectTTS w/o sub-match (inv k-means (avg)) 3.89 $\pm$ 0.08 3.71 $\pm$ 0.09
SelectTTS w sub-match (inv k-means (avg)) 3.97 $\pm$ 0.08 3.72 $\pm$ 0.09
@@ V-B Subjective evaluation
We conduct listening experiments with 21 participants to evaluate the naturalness and speaker similarity of the synthesized speech. A total of 120 utterances from 6 unseen speakers (3 male, 3 female) from the LibriTTS-R test set are used for the evaluation. In the Mean Opinion Score (MOS) test [7], participants rate the naturalness of speech samples on a 5-point scale, evaluating 12 utterances per method. As shown in Table V, SelectTTS achieves a high MOS score close to 4, comparable to strong baselines. Sub-sequence matching further improves the perceived naturalness. We also conduct a Speaker Mean Opinion Score (SMOS) evaluation [7], where participants are presented with a ground-truth recording and a synthesized utterance and asked to rate speaker similarity on a 5-point scale. SelectTTS achieves strong SMOS results, demonstrating speaker resemblance on par with or better than larger baseline models.
@@ VI Conclusion
In this paper, we propose SelectTTS, a multi-speaker TTS framework with lower model complexity that synthesizes high-quality speech closely resembling the target speaker’s voice by directly utilizing frames from unseen speakers. We demonstrate that combining semantic unit-based frame selection with vocoding over SSL feature sequences provides a simple yet effective approach for modeling unseen speakers, achieving state-of-the-art speaker similarity. SelectTTS significantly reduces model complexity and data requirements, opening possibilities for extending state-of-the-art multi-speaker TTS to low-resource settings and languages.